OpenStudy (anonymous):

solve for x (4-x)(5-x)=3

OpenStudy (hoblos):

first expand (4-x)(5-x) = 20 -4x -5x + x^2 = x^2-9x+20 so x^2-9x+20=3 x^2 - 9x + 17 = 0 it is a quadratic now... do you know how to solve it?

OpenStudy (anonymous):

6.302775637, 2.697224362 is this right?

OpenStudy (hoblos):

using the calculayor... yes you are right but dont you know the quadratic formula to find the solutions?

OpenStudy (anonymous):

yeah -b +or- squareroot of b*2-4ac / 2a

OpenStudy (anonymous):

or is that wrong?

OpenStudy (hoblos):

yes thats it.. and its better to use it.. it gives the exact values